Maria practiced the recorder for 40 minutes on Saturday. She practiced 10 minutes less in the afternoon than in the mornning. how many minutes did maria practice in the morning? in the afternoon?

Answer by fractalier(2101)

Let M be her morning time and A be her afternoon time.
So we know that
M + A = 40
and
A = M - 10
We can substitute the second into the first and get
M + (M - 10) = 40
2M - 10 = 40
2M = 50
M = 25 min
and then
A = 15 min